# Shifting AI Efficiency From Model-Centric to Data-Centric Compression

### v1

@@ 1 Introduction
The explosive growth of large language models (LLMs) (OpenAI:GPT-4, ; Touvron:LLaMA, ; Bai:Qwen, ; grattafiori2024llama3, ; dong2024internlm, ; yang2024Qwen2, ; yang2025Qwen3, ; liu2024deepseek-v3, ; guo2025deepseek-r1, ) and their multi-modal extensions (MLLMs) (liu2023llava, ; Liu:LLaVA-1.5, ; Chen:InternVL, ; chen2024internvl-2.5, ; zhu2025internvl3, ; wang2024qwen2-vl, ; bai2025qwen2.5-vl, ; chu2024qwen2-audio, ) over the past few years has driven remarkable gains in AI capabilities. Notably, this unprecedented progress has been largely achieved through increasing model scale across the field, with larger models consistently demonstrating superior performance in reasoning, knowledge acquisition, and task generalization. Indeed, the evolution from early language models with modest parameter counts, such as BERT (117M) (devlin2018bert, ), to today’s state-of-the-art LLMs like Llama 4 (llama4, ), DeepSeek-R1 (guo2025deepseek-r1, ), and Qwen-3 (yang2025Qwen3, ) (100B+), demonstrates how each successive model iteration has delivered disproportionate performance improvements through sheer scale. Nevertheless, this relentless pursuit of enhanced performance through increased model size comes at an ever-increasing computational cost. As a result, by early 2024, the dominant source of computational overhead was primarily attributed to the linear growth in parameter count and associated memory requirements.
@@
In response to this scaling trend, the research community has developed numerous model-centric compression techniques. These include model quantization (yang2019quantization, ; rokh2023comprehensive, ), network pruning (Deep_Compression, ; cheng2024survey, ), knowledge distillation (Hinton:KD, ; gou2021KD_Survey, ) and low-rank decomposition yu2017low-rank ; idelbayev2020low-rank . These methods directly reduce computational overhead by decreasing model size, and were a natural response to the 2022 to 2024 era, when scaling up model size was the primary driver of performance gains.
@@
As model sizes approach hardware limits, the pace of parameter growth is flattening. Meanwhile, a new computational challenge has emerged: the exponential growth in token sequence lengths. Figure 1 (left) clearly shows that while from 2022 to 2024, model size primarily drove computational costs, reaching around 1000B parameters before stagnating, from 2024 onward, the dominant factor has dramatically shifted to the staggering number of processed tokens, which continues to grow exponentially. This trend is evident across multiple domains: language models now process context lengths orders of magnitude longer than their predecessors (llama4, ; yang2025Qwen3, ; liu2024deepseek-v3, ), particularly with emerging technologies like long chain-of-thought reasoning (guo2025deepseek-r1, ) and multi-agent systems (han2024llm-multi-agent, ), vision models must handle increasingly high-resolution images (li2024llava-ov, ; zhu2025internvl3, ; bai2025qwen2.5-vl, ) and longer videos (chen2024longvila, ; liu2025video-xl-pro, ), and generative models are tasked with creating higher-resolution images (chen2024pixart-sigma, ; flux2024, ) and hour-long videos (openai2024sora, ), requiring substantially more tokens and leading to overwhelming computational overhead. Consequently, by late 2024, the primary computational bottleneck has clearly shifted from model size to the quadratic cost of attention mechanism over these extremely long token sequences.
@@
This unprecedented growth in sequence lengths has fundamentally shifted the computational bottleneck from model size to the quadratic cost of attention over long context sequences. Based on this observation, as illustrated in Figure 1 (right), we propose a critical position: the AI community should shift its efficiency optimization paradigm from model-centric to data-centric compression. Specifically, we advocate for token compression - a data-centric compression approach that directly reduces token redundancy in model inputs (goyal2020power, ; jiang2023llmlingua, ; li2024snapkv, ; Bolya2023:ToMe, ; Bolya2023:ToMeSD, ; Chen2024:FastV, ). Token compression methods address computational overhead by identifying and removing low-information tokens during processing, generally without modifying model architectures or even requiring retraining. Our detailed analysis in Section 3.3 demonstrates that token compression offers compelling advantages in terms of universality, efficiency, and compatibility, positioning it as a promising solution for enabling efficient next-generation LLMs and MLLMs.
@@ 2 Background
2.1 Token Overhead aross Various Domains
The field of artificial intelligence has witnessed remarkable advancements across multiple domains, including natural language processing, computer vision, and content generation. These developments have been largely driven by the introduction of the Transformer architecture (vaswani2017:Transformer, ), which has spawned a wide variety of models. As these domains evolve, we observe a significant increase in token sequence lengths across three main areas:
$\bullet$ Longer Context Length in Language Models Large language models (LLMs) (devlin2018bert, ; Touvron:LLaMA, ; grattafiori2024llama3, ; liu2024deepseek-v3, ; Bai:Qwen, ; yang2024Qwen2, ; yang2025Qwen3, ; OpenAI:GPT-4, ; cai2024internlm2, ) have demonstrated remarkable capabilities in natural language understanding and generation (zhao2023survey, ; hadi2023survey, ). The context length these models can handle has expanded dramatically from 2,048 tokens in early models like Llama 1 (Touvron:LLaMA, ) to 10M tokens in recent iterations like Llama 4 Scout (llama4, ). This expansion has led to the emergence of large reasoning models (guo2025deepseek-r1, ; hurst2024gpt, ; huang2022towards, ), which focus on complex multi-step problem solving through techniques like long chain-of-thought reasoning (liu2024deepseek, ; yang2025Qwen3, ; chen2025longcot, ) and multi-agent collaboration (talebirad2023Multi-agent-collaboration, ; chen2024pandora, ; han2024llm-multi-agent, ).
$\bullet$ Higher Resolution and Longer Video Understanding Building on the success of LLMs, multi-modal large language models (MLLMs) (li2022blip, ; Liu:LLaVA-1.5, ; Liu:LLaVA-NeXT, ; li2024llava-ov, ; Bai:Qwen-VL, ; Chen:InternVL, ; chen2024internvl-2.5, ; zhu2025internvl3, ) extend these capabilities by integrating vision and text processing (wu2023multimodal, ). The visual inputs these models process have evolved significantly from basic $224\times 224$ resolution images in early models like LLaVA (liu2023llava, ) to 4K ultra-high-resolution images in InternVL3 (zhu2025internvl3, ) and 8K-frame videos in Video-XL-Pro (liu2025video-xl-pro, ), achieving remarkable performance in tasks involving images (bai2025qwen2.5-vl, ), videos (chen2024longvila, ), and multi-modal reasoning (wang2024exploring, ).
$\bullet$ More Complex Content in Generation Tasks The field of content generation has seen dramatic advancements, particularly with the application of Transformers to generative domains (peebles2023DiT, ; openai2024sora, ; li2024hunyuandit, ; chen2024mj, ). Early diffusion models like Stable Diffusion (rombach2022SD, ) were limited to generating $512\times 512$ resolution images. With Transformers being successfully applied to generative domains (peebles2023DiT, ; gao2023maskdit, ; openai2024sora, ; li2024hunyuandit, ), DiT-based models have dramatically advanced the field, producing high-quality 4K images in PixArt-$\Sigma$ (chen2024pixart-sigma, ) and even hour-long videos in Sora (openai2024sora, ). These models capture complex dependencies across space and time, achieving remarkable results in high-fidelity content generation (flux2024, ; yang2024cogvideox, ; wan2025wan, ; kang2025legion, ).
@@ 2.2 Model Efficiency from Different Perspectives
where model efficiency can be optimized from three perspectives: (I) Efficient Computation Architecture aims to design efficient neural architectures $\mathbf{F}$ (shen2021-liner-attention, ; peng2023rwkv, ; gu2023mamba, ), (II) Model-centric Compression focuses on model weights $\mathbf{W}$ (Hinton:KD, ; yang2019quantization, ; li2017pruning, ; yu2017low-rank, ), and (III) Data-centric Compression targets token sequences from input data $\mathbf{X}$ (Rao2021:DynamicViT, ; jiang2023llmlingua, ; Bolya2023:ToMe, ; zou2025:ToCa, ; Chen2024:FastV, ).
@@
Since the computational efficiency of neural networks is determined by their architectural design, optimizing architectures represents a fundamental approach to enhance efficiency. Unlike traditional Transformer architectures with quadratic computational complexity $\mathcal{O}(n^{2})$ in attention (vaswani2017:Transformer, ), where $n$ is sequence length, recent innovations typically achieve linear or sub-quadratic scaling: (i) linear attention reformulates the attention mechanism to achieve linear complexity $\mathcal{O}(n)$, enabling efficient processing while maintaining model capacity (katharopoulos2020:transformers-rnn, ; shen2021-liner-attention, ); (ii) RWKV architecture integrates RNN-like linear scaling $\mathcal{O}(n)$ with transformer-like parallelism for efficient sequence processing (peng2023rwkv, ; duan2024visionrwkv, ); (iii) State Space Models like Mamba leverage structured state space modeling to achieve linear complexity $\mathcal{O}(n)$ while maintaining performance (gu2023mamba, ; Zhu2024:Vim, ).While these architectural innovations improve efficiency significantly, they require complete model retraining, motivating the exploration of alternative approaches.
@@
where $\boldsymbol{\Gamma}$ represents the compression operator and ||$\cdot$|| ⋅ | denotes parameter size. Several key approaches have emerged: (i) network pruning removes redundant weights, reducing parameters and computation (Network_Pruning, ; cheng2024survey, ); (ii) model quantization reduces parameter precision from high-bit to low-bit representations (yang2019quantization, ; rokh2023comprehensive, ); (iii) knowledge distillation transfers information from large models to compact ones Hinton:KD ; gou2021KD_Survey ; (iv) low-rank decomposition approximates weight matrices with lower-rank representations (yu2017low-rank, ; idelbayev2020low-rank, ). With model sizes plateauing and context lengths growing, research focus has begun shifting from model-level to token-level compression strategies.
@@
where $\boldsymbol{\Phi}$ is the token compression operator and ||$\cdot$|| ⋅ | denotes token length. This approach complements model compression and has demonstrated significant effectiveness across computer vision (Rao2021:DynamicViT, ; Bolya2023:ToMe, ) and natural language processing (kim2021length, ; jiang2023llmlingua, ).
@@ 3.1 Research Roadmap - What Makes Token Compression Work?
(I) Parametric Methods employ auxiliary networks as scoring functions $\mathcal{E}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, introducing additional parameters $\Delta\theta$ beyond the original model parameters $\theta$. These methods include: (i) training-aware approaches (Rao2021:DynamicViT, ; you2024DiffCR, ; li2024tokenpacker, ; kim2022learned, ) that optimize $\Delta\theta$ through training to learn scoring function $\mathcal{S}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, and (ii) training-free approaches (Mahmud2024:PaPr, ; zhao2024SGL, ) that directly employ pre-trained networks as scoring function $\mathcal{S}_{\text{fixed}}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$ without updating $\Delta\theta$.
(II) Non-parametric Methods utilize parameter-free heuristics for token scoring without introducing extra parameters. These approaches can be categorized into: (i) inherent computation methods (Liang2022:EViT, ; zou2025:ToCa, ; Chen2024:FastV, ; xiao2024efficient, ; ge2024model, ) that leverage model’s internal calculations for token scoring $\mathcal{S}_{\text{in}}:\mathbf{A}\to\{s_{t}\}_{t=1}^{T}$, such as using attention weights ($s_{t}=\sum_{j=1}^{T}a_{t}^{j}$, where $a_{t}^{j}$ represents attention score between tokens), and (ii) external computation methods (bolya2023tome, ; Zhang2025:SiTo, ; devoto-etal-2024-simple, ; wang2024model, ; liu2025vidcom2, ) that design additional metrics $\mathcal{S}_{\text{ex}}:\mathbf{Z}\to\mathbb{R}^{T\times T}$ to evaluate token relationships. For external methods, an additional function $g:\mathbf{X}\to\mathbf{Z}$ is introduced to compute intermediate features, where $\mathbf{Z}=g(\mathbf{X})$. The scoring function then operates on these features: $s_{i,j}=f(\mathbf{z}_{i},\mathbf{z}_{j})$, where $f$ is a custom pairwise scoring function. A typical example is using cosine similarity, where $g$ is an identity function and $s_{i,j}=\frac{\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle}{\|\mathbf{x}_{i}\|_{2}\|\mathbf{x}_{j}\|_{2}}$.
@@
(I) Token Pruning directly discards less important tokens from the sequence based on their scores. These methods (Rao2021:DynamicViT, ; goyal2020power, ; jiang2023llmlingua, ; Chen2024:FastV, ) typically remove tokens with scores below a threshold, producing a compressed sequence:
@@
(II) Token Merging preserves information by combining semantically similar tokens (bolya2023tome, ; Liang2022:EViT, ; zhang2024sparsevlm, ; Bolya2023:ToMeSD, ). Given an input sequence $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ and a mapping $\pi:\{1,\dots,T\}\to\{1,\dots,M\}$ that assigns tokens to $M$ merge groups based on their semantic relationships, this approach generates a compressed sequence $\mathbf{X}^{\prime}=\{\mathbf{x}^{\prime}_{1},\dots,\mathbf{x}^{\prime}_{M}\}$ through weighted aggregation:

### v2

@@ 1 Introduction
The explosive growth of large language models (LLMs) OpenAI:GPT-4 ; Touvron:LLaMA ; Bai:Qwen ; grattafiori2024llama3 ; dong2024internlm ; yang2024Qwen2 ; yang2025Qwen3 ; liu2024deepseek-v3 ; guo2025deepseek-r1 and their multi-modal extensions (MLLMs) liu2023llava ; Liu:LLaVA-1.5 ; Chen:InternVL ; chen2024internvl-2.5 ; zhu2025internvl3 ; wang2024qwen2-vl ; bai2025qwen2.5-vl ; chu2024qwen2-audio over the past few years has driven remarkable gains in AI capabilities. Notably, this unprecedented progress has been largely achieved through increasing model scale across the field, with larger models consistently demonstrating superior performance in reasoning, knowledge acquisition, and task generalization. Indeed, the evolution from early language models with modest parameter counts, such as BERT (117M) devlin2018bert , to today’s state-of-the-art LLMs like Llama 4 llama4 , DeepSeek-R1 guo2025deepseek-r1 , and Qwen-3 yang2025Qwen3 (100B+), demonstrates how each successive model iteration has delivered disproportionate performance improvements through sheer scale. Nevertheless, this relentless pursuit of enhanced performance through increased model size comes at an ever-increasing computational cost. As a result, by early 2024, the dominant source of computational overhead was primarily attributed to the linear growth in parameter count and associated memory requirements.
@@
In response to this scaling trend, the research community has developed numerous model-centric compression techniques. These include model quantization yang2019quantization ; rokh2023comprehensive , network pruning Deep_Compression ; cheng2024survey , knowledge distillation Hinton:KD ; gou2021KD_Survey and low-rank decomposition yu2017low-rank ; idelbayev2020low-rank . These methods directly reduce computational overhead by decreasing model size, and were a natural response to the 2022 to 2024 era, when scaling up model size was the primary driver of performance gains.
@@
As model sizes approach hardware limits, the pace of parameter growth is flattening. Meanwhile, a new computational challenge has emerged: the exponential growth in token sequence lengths. Figure 1 (left) clearly shows that while from 2022 to 2024, model size primarily drove computational costs, reaching around 1000B parameters before stagnating, from 2024 onward, the dominant factor has dramatically shifted to the staggering number of processed tokens, which continues to grow exponentially. This trend is evident across multiple domains: language models now process context lengths orders of magnitude longer than their predecessors llama4 ; yang2025Qwen3 ; liu2024deepseek-v3 , particularly with emerging technologies like long chain-of-thought reasoning guo2025deepseek-r1 and multi-agent systems han2024llm-multi-agent , vision models must handle increasingly high-resolution images li2024llava-ov ; zhu2025internvl3 ; bai2025qwen2.5-vl and longer videos chen2024longvila ; liu2025video-xl-pro , and generative models are tasked with creating higher-resolution images chen2024pixart-sigma ; flux2024 and hour-long videos openai2024sora , requiring substantially more tokens and leading to overwhelming computational overhead. Consequently, by late 2024, the primary computational bottleneck has clearly shifted from model size to the quadratic cost of attention mechanism over these extremely long token sequences.
@@
This unprecedented growth in sequence lengths has fundamentally shifted the computational bottleneck from model size to the quadratic cost of attention over long context sequences. Based on this observation, as illustrated in Figure 1 (right), we propose a critical position: the AI community should shift its efficiency optimization paradigm from model-centric to data-centric compression. Specifically, we advocate for token compression - a data-centric compression approach that directly reduces token redundancy in model inputs goyal2020power ; jiang2023llmlingua ; li2024snapkv ; Bolya2023:ToMe ; Bolya2023:ToMeSD ; Chen2024:FastV . Token compression methods address computational overhead by identifying and removing low-information tokens during processing, generally without modifying model architectures or even requiring retraining. Our detailed analysis in Section 3.3 demonstrates that token compression offers compelling advantages in terms of universality, efficiency, and compatibility, positioning it as a promising solution for enabling efficient next-generation LLMs and MLLMs.
@@ 2 Background
2.1 Token Overhead across Various Domains
The field of artificial intelligence has witnessed remarkable advancements across multiple domains, including natural language processing, computer vision, and content generation. These developments have been largely driven by the introduction of the Transformer architecture vaswani2017:Transformer , which has spawned a wide variety of models. As these domains evolve, we observe a significant increase in token sequence lengths across three main areas:
$\bullet$ Longer Context Length in Language Models Large language models (LLMs) devlin2018bert ; Touvron:LLaMA ; grattafiori2024llama3 ; liu2024deepseek-v3 ; Bai:Qwen ; yang2024Qwen2 ; yang2025Qwen3 ; OpenAI:GPT-4 ; cai2024internlm2 have demonstrated remarkable capabilities in natural language understanding and generation zhao2023survey ; hadi2023survey . The context length these models can handle has expanded dramatically from 2,048 tokens in early models like Llama 1 Touvron:LLaMA to 10M tokens in recent iterations like Llama 4 Scout llama4 . This expansion has led to the emergence of large reasoning models guo2025deepseek-r1 ; hurst2024gpt ; huang2022towards , which focus on complex multi-step problem solving through techniques like long chain-of-thought reasoning liu2024deepseek ; yang2025Qwen3 ; chen2025longcot and multi-agent collaboration talebirad2023Multi-agent-collaboration ; chen2024pandora ; han2024llm-multi-agent .
$\bullet$ Higher Resolution and Longer Video Understanding Building on the success of LLMs, multi-modal large language models (MLLMs) li2022blip ; Liu:LLaVA-1.5 ; Liu:LLaVA-NeXT ; li2024llava-ov ; Bai:Qwen-VL ; Chen:InternVL ; chen2024internvl-2.5 ; zhu2025internvl3 extend these capabilities by integrating vision and text processing wu2023multimodal . The visual inputs these models process have evolved significantly from basic $\times 224$ 224 resolution images in early models like LLaVA liu2023llava to 4K ultra-high-resolution images in InternVL3 zhu2025internvl3 and 8K-frame videos in Video-XL-Pro liu2025video-xl-pro , achieving remarkable performance in tasks involving images bai2025qwen2.5-vl , videos chen2024longvila , and multi-modal reasoning wang2024exploring .
[…]
@@ 2.2 Model Efficiency from Different Perspectives
where model efficiency can be optimized from three perspectives: (I) Efficient Computation Architecture aims to design efficient neural architectures $\mathbf{F}$ shen2021-liner-attention ; peng2023rwkv ; gu2023mamba , (II) Model-centric Compression focuses on model weights $\mathbf{W}$ Hinton:KD ; yang2019quantization ; li2017pruning ; yu2017low-rank , and (III) Data-centric Compression targets token sequences from input data $\mathbf{X}$ Rao2021:DynamicViT ; jiang2023llmlingua ; Bolya2023:ToMe ; zou2025:ToCa ; Chen2024:FastV .
@@
Since the computational efficiency of neural networks is determined by their architectural design, optimizing architectures represents a fundamental approach to enhance efficiency. Unlike traditional Transformer architectures with quadratic computational complexity $\mathcal{O}(n^{2})$ in attention vaswani2017:Transformer , where $n$ is sequence length, recent innovations typically achieve linear or sub-quadratic scaling: (i) linear attention reformulates the attention mechanism to achieve linear complexity $\mathcal{O}(n)$, enabling efficient processing while maintaining model capacity katharopoulos2020:transformers-rnn ; shen2021-liner-attention ; (ii) RWKV architecture integrates RNN-like linear scaling $\mathcal{O}(n)$ with transformer-like parallelism for efficient sequence processing peng2023rwkv ; duan2024visionrwkv ; (iii) State Space Models like Mamba leverage structured state space modeling to achieve linear complexity $\mathcal{O}(n)$ while maintaining performance gu2023mamba ; Zhu2024:Vim .While these architectural innovations improve efficiency significantly, they require complete model retraining, motivating the exploration of alternative approaches.
@@
where $\boldsymbol{\Gamma}$ represents the compression operator and ||$\cdot$|| ⋅ | denotes parameter size. Several key approaches have emerged: (i) network pruning removes redundant weights, reducing parameters and computation Network_Pruning ; cheng2024survey ; (ii) model quantization reduces parameter precision from high-bit to low-bit representations yang2019quantization ; rokh2023comprehensive ; (iii) knowledge distillation transfers information from large models to compact ones Hinton:KD ; gou2021KD_Survey ; (iv) low-rank decomposition approximates weight matrices with lower-rank representations yu2017low-rank ; idelbayev2020low-rank . With model sizes plateauing and context lengths growing, research focus has begun shifting from model-level to token-level compression strategies.
@@
where $\boldsymbol{\Phi}$ is the token compression operator and ||$\cdot$|| ⋅ | denotes token length. This approach complements model compression and has demonstrated significant effectiveness across computer vision Rao2021:DynamicViT ; Bolya2023:ToMe and natural language processing kim2021length ; jiang2023llmlingua .
@@ 3.1 Research Roadmap - What Makes Token Compression Work?
(I) Parametric Methods employ auxiliary networks as scoring functions $\mathcal{E}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, introducing additional parameters $\Delta\theta$ beyond the original model parameters $\theta$. These methods include: (i) training-aware approaches Rao2021:DynamicViT ; you2024DiffCR ; li2024tokenpacker ; kim2022learned that optimize $\Delta\theta$ through training to learn scoring function $\mathcal{S}_{\Delta\theta}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$, and (ii) training-free approaches Mahmud2024:PaPr ; zhao2024SGL that directly employ pre-trained networks as scoring function $\mathcal{S}_{\text{fixed}}:\mathbf{X}\to\{s_{t}\}_{t=1}^{T}$ without updating $\Delta\theta$.
(II) Non-parametric Methods utilize parameter-free heuristics for token scoring without introducing extra parameters. These approaches can be categorized into: (i) inherent computation methods Liang2022:EViT ; zou2025:ToCa ; Chen2024:FastV ; xiao2024efficient ; ge2024model that leverage model’s internal calculations for token scoring $\mathcal{S}_{\text{in}}:\mathbf{A}\to\{s_{t}\}_{t=1}^{T}$, such as using attention weights ($s_{t}=\sum_{j=1}^{T}a_{t}^{j}$, where $a_{t}^{j}$ represents attention score between tokens), and (ii) external computation methods bolya2023tome ; Zhang2025:SiTo ; devoto-etal-2024-simple ; wang2024model ; liu2025vidcom2 that design additional metrics $\mathcal{S}_{\text{ex}}:\mathbf{Z}\to\mathbb{R}^{T\times T}$ to evaluate token relationships. For external methods, an additional function $g:\mathbf{X}\to\mathbf{Z}$ is introduced to compute intermediate features, where $\mathbf{Z}=g(\mathbf{X})$. The scoring function then operates on these features: $s_{i,j}=f(\mathbf{z}_{i},\mathbf{z}_{j})$, where $f$ is a custom pairwise scoring function. A typical example is using cosine similarity, where $g$ is an identity function and $s_{i,j}=\frac{\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle}{\|\mathbf{x}_{i}\|_{2}\|\mathbf{x}_{j}\|_{2}}$.
@@
(I) Token Pruning directly discards less important tokens from the sequence based on their scores. These methods Rao2021:DynamicViT ; goyal2020power ; jiang2023llmlingua ; Chen2024:FastV typically remove tokens with scores below a threshold, producing a compressed sequence:
@@
(II) Token Merging preserves information by combining semantically similar tokens bolya2023tome ; Liang2022:EViT ; zhang2024sparsevlm ; Bolya2023:ToMeSD . Given an input sequence $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ and a mapping $\pi:\{1,\dots,T\}\to\{1,\dots,M\}$ that assigns tokens to $M$ merge groups based on their semantic relationships, this approach generates a compressed sequence $\mathbf{X}^{\prime}=\{\mathbf{x}^{\prime}_{1},\dots,\mathbf{x}^{\prime}_{M}\}$ through weighted aggregation: